OpenStudy (anonymous):

Given that f(x) = –x + 4 and g(x) = –2x – 3, solve for f(g(x)) when x = 2.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf f(x)=-x+4\qquad {\color{brown}{ g(x)}}=-2x-3 \\ \quad \\ f(\quad {\color{brown}{ g(x)}}\quad )=-{\color{brown}{ g(x)}}+4\implies f(\quad {\color{brown}{ g(x)}}\quad )=-{\color{brown}{ (-2x-3)}}+4 \\ \quad \\ f(\quad {\color{brown}{ g({\color{blue}{ 2}})}}\quad )=-{\color{brown}{ (-2({\color{blue}{ 2}})-3)}}+4\)

OpenStudy (anonymous):

I'm still lost... this is my first day of Algebra 2.

OpenStudy (jdoe0001):

f( g(x) ) means, "replace any "x" variable in f(x), with g(x)" notice how the g(x) takes over the "x" in f(x)

OpenStudy (jdoe0001):

for example if say g(x) = ducks then f( g(x) ) = -(ducks) + 4 ^ used to be x

OpenStudy (anonymous):

So do I just plug in 2 somewhere?

OpenStudy (jdoe0001):

well... first you expand the f( g(x) ) and then plug in the "2"

OpenStudy (anonymous):

Ahhh I'm really struggling here.

OpenStudy (jdoe0001):

f( something ) = -"something" + 4 f( pelicans ) = -"pelicans" + 4 f( albatross ) = - "albatross" + 4 "x" is just a variable, it coiuld be anything, it just so happen that's "x" but whatever f( "show here" ) <-- will be the variable used in it

OpenStudy (anonymous):

Ohh.

OpenStudy (jdoe0001):

say for example \(\bf f(x) = 2x+3x^2\qquad g(x)=pelicans \\ \quad \\ f(\quad g(x)\quad ) \iff f(pelicans)=2(pelicans)+3(pelicans)^2\)

OpenStudy (anonymous):

Ah okay!

OpenStudy (anonymous):

(Still a bit confused but I'm starting to get it)

OpenStudy (anonymous):

How would this information help me create a solution?

OpenStudy (jdoe0001):

well... notice how g(x) gets inserted firstly and then how the "2" gets inserted secondly \(\bf f(x)=-x+4\qquad {\color{brown}{ g(x)}}=-2x-3 \\ \quad \\ f(\quad {\color{brown}{ g(x)}}\quad )=-{\color{brown}{ g(x)}}+4\implies f(\quad {\color{brown}{ g(x)}}\quad )=-{\color{brown}{ (-2x-3)}}+4 \\ \quad \\ f(\quad {\color{brown}{ g({\color{blue}{ 2}})}}\quad )=-{\color{brown}{ (-2({\color{blue}{ 2}})-3)}}+4\)

OpenStudy (anonymous):

I got 11.

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

..OH!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

Would you mind doing one more? It's a fraction one.

OpenStudy (jdoe0001):

easier if you post anew, sure... that way if I dunno, someone else may

OpenStudy (anonymous):

Posted!